OpenStudy (anonymous):

By what approximate factor is the intensity of an earthquake with magnitude 5.4 greater than an earthquake with magnitude 5.3? M = the magnitude of an earthquake I = the intensity of an earthquake I0 = the smallest seismic activity that can be measured

OpenStudy (anonymous):

M = log (I/I0)

OpenStudy (anonymous):

@MrCoolGuy

OpenStudy (anonymous):

@haleyelizabeth2017

OpenStudy (anonymous):

@Qwertty123

OpenStudy (anonymous):

Will medal and fan

OpenStudy (qwertty123):

okay so what do you have so far?

OpenStudy (anonymous):

1.01 1.21 1.26 10.44

OpenStudy (anonymous):

i would say 1.01

OpenStudy (anonymous):

due to 5.3 and 5.4 are only .1 apart

OpenStudy (anonymous):

am i right? @Qwertty123

OpenStudy (qwertty123):

Not sure.. @mathstudent55

OpenStudy (anonymous):

How does the value of a affect the graph of a monomial function?

OpenStudy (anonymous):

As |a| increases, the graph widens; as |a| decreases, the graph narrows. As |a| increases, the graph narrows; as |a| decreases, the graph widens. As |a| increases, the graph translates to the right; as |a| decreases, the graph translates to the left. As |a| increases, the graph translates up; as |a| decreases, the graph translates down.

OpenStudy (anonymous):

what about this one?

OpenStudy (anonymous):

@Qwertty123

OpenStudy (qwertty123):

not sure.. @phi

OpenStudy (phi):

is it \[ M = \log \left(\frac{I}{I_0}\right)\] ? if so \[ 10^M = \frac{I}{I_0} \\ I = I_0 10^M\] you want the "factor" (call it "a") \[ I_{5.4} = a I_{5.3} \] or replacing the I's \[ I_0 10^{5.4} = a \ I_0 10^{5.3} \] now solve for "a"

OpenStudy (phi):

you get \[ a = \frac{10^{5.4}}{10^{5.3}} \] you can simplify because (with the same base) we can subtract exponents: \[ a = 10^{0.1} \]